A flywheel is rotating at 1200 rpm when a brake is applied. If it comes to rest in 13.4 seconds, what is the angular acceleration in rad/s2? Be careful with signs! Express your result to three significant digits.
Answer #1

w = 1200 rpm = 1200*2\pi/60 = 125.6 rad /s

angular acceleration \alpha = (w-wo)/t = (0-125.6 )/13.4

= - 9.37 rad/s2